OpenStudy (anonymous):

how would you solve this equation: 7x-1=3y+8

OpenStudy (anonymous):

In slope intercept form?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

use the same technique as your other equations. You can start by subtracting 8 from both sides. I will help you along the way, but try to work through it.

OpenStudy (anonymous):

ok so 7x-1=3y+8 -8 -8 -1x-1=3y

OpenStudy (anonymous):

8 is not multiplied by x, so you cannot subtract it from 7x. It would instead be 7x-1-8=7x-9. Make sense?

OpenStudy (anonymous):

oh ok understand now thank you!

OpenStudy (anonymous):

No problem. Post your answer if you want confirmation.

OpenStudy (anonymous):

is the answer: y=7x-3? or y=7x-9?

OpenStudy (anonymous):

The first one is closest, but you forgot to divide all parts by 3... 7x-9=3y You have to divide all of it: (7x-9)/3

OpenStudy (anonymous):

so y=7/3x-3?

OpenStudy (anonymous):

Yep - that's it!